# Max–Min Fairness in Stacked Intelligent Metasurface-Aided Rate Splitting Networks Abdullah Quran, Shimaa Naser, Maryam Tariq, Omar Alhussein, and Sami Muhaidat

### v1

@@ III Joint Power Allocation and Wave-Based Beamforming: Solution to Problem (P1)
To handle the non-convexity of Problem (P1), we adopt an AO approach that alternately optimizes the power allocation and SIM phase shift variables. The corresponding subproblems are introduced in the following subsections.
@@ III-A Power Allocation
[…]
where $t$ is an auxiliary variable introduced to represent the minimum rate across all users and to enable a smooth reformulation of the original max–min objective. However, problem (P2) remains non-convex due to the non-convexity of constraint (5d) and constraint (6b), since both the private rates $R_{p,k}$ and the common rates $R_{c,k}$ are non-concave functions of the power allocation vector $\mathbf{p}$.
@@
$\footnotesize R_{c,k}=\log_{2}(p_{c}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{c}|^{2}+\sum_{j\in\mathcal{K}}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2})-\eta_{c,k},$ (7)
$\footnotesize R_{p,k}=\log_{2}(\sum_{j\in\mathcal{K}}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2})-\eta_{p,k},$ (8)
@@
$\footnotesize\eta_{\delta,k}=\log_{2}\left(\sum_{j\in\mathcal{K}_{\delta}(k)}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2}\right)\!,\quad\delta\in\{\mathrm{c},\mathrm{p}\}.$ (9)
Despite this reformulation, $R_{c,k}$ and $R_{p,k}$ remain non-concave due to the non-concavity of the term $\eta_{\delta,k}$. To address this, we approximate (9) using first-order Taylor expansions around the solution obtained at iteration $t-1$ as follows
$\footnotesize\tilde{\eta}_{\delta,k}^{(t)}\!=\!\eta_{\delta,k}^{(t-1)}+\frac{\sum_{j\in\mathcal{K}_{\delta}(k)}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}(p_{j}-p_{j}^{(t-1)})}{\ln 2\left(\sum_{i\in\mathcal{K}_{\delta}(k)}p_{i}^{(t-1)}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{i}|^{2}+\sigma_{k}^{2}\right)}.$ (10)
Finally, $R_{c,k}$ and $R_{p,k}$ are now concave functions of the power allocation vector $\mathbf{p}$. Consequently, constraints (5d) and (6b) are convex, and the approximate problem at each iteration is convex and can be efficiently solved using convex optimization solvers such as CVX [17]. Hence, we employ the SCA method, which iteratively refines the solution by solving a sequence of convex subproblems. At iteration $t$, SCA approximates the non-concave components $\eta_{c,k}$ and $\eta_{p,k}$ with their first-order convex surrogates $\tilde{\eta}_{c,k}^{(t)}$ and $\tilde{\eta}_{p,k}^{(t)}$, evaluated using the solution from the previous iteration following (10).
@@ III-B SIM Beamforming Optimization
Given a fixed power allocation, the SIM beamforming subproblem aims to maximize the minimum rate across all users by optimizing the phase shifts applied across all metasurface layers. The SIM beamforming subproblem is expressed as
[…]
To facilitate the solution to (P3), we collect the diagonal elements of the phase shift matrices $\{\boldsymbol{\Theta}_{\ell}\}_{\ell=1}^{L}$ into a single stacked vector $\boldsymbol{\theta}\in\mathbb{C}^{LM\times 1}$, defined as $\boldsymbol{\theta}=[\theta_{1,1},\dots,\theta_{1,M},\theta_{2,1},\dots,\theta_{L,M}]^{T}$, where each element $\theta_{\ell,m}$ satisfies the unit-modulus constraint (5b). The non-convex nature of this constraint makes it particularly difficult to handle using traditional optimization techniques, necessitating more advanced solution methods.
To handle the non-convexity of constraint (5b), we reinterpret the feasible set defined by (5b) as a Riemannian manifold[18] $\mathcal{S}$ as follows
$\footnotesize\mathcal{S}=\{\boldsymbol{\theta}\in\mathbb{C}^{LM}:|\theta_{i}|=1,\quad\forall i\in\{1,\dots,LM\},$ (12)
@@
Moreover, to address the non-smoothness of the objective function in (11a), we adopt a log-sum-exp surrogate to obtain a differentiable approximation:
$\footnotesize\mathcal{F}(\boldsymbol{\theta})=-\frac{1}{\beta}\log\left(\sum_{k\in\mathcal{K}}e^{-\beta(R_{c,k}+R_{p,k})}\right)\!,$ (13)
where $\beta>0$ is a parameter that controls the approximation tightness. With this smooth reformulation, we develop a Riemannian Conjugate Gradient (RCG) method over the manifold $\mathcal{S}$ to optimize the SIM phase shifts. In each iteration of this method, we compute the Euclidean gradient, $\nabla_{\boldsymbol{\theta}}\mathcal{F}$. Because updates based on the Euclidean gradient may not respect the manifold constraint $\mathcal{S}$, we project it onto the tangent space $T_{\boldsymbol{\theta}}\mathcal{S}$ to obtain the Riemannian gradient:
$\footnotesize\mathbf{g}_{\boldsymbol{\theta}}=\nabla_{\boldsymbol{\theta}}\mathcal{F}-\Re\left(\nabla_{\boldsymbol{\theta}}\mathcal{F}\odot\boldsymbol{\theta}^{*}\right)\odot\boldsymbol{\theta},$ (14)
@@
$\footnotesize\mathbf{u}^{(t+1)}=\mathbf{g}^{(t+1)}+\gamma^{(t+1)}\mathbf{u}^{(t)},$ (4)
where $\gamma^{(t+1)}$ is computed using the Polak–Ribière formula [18]. Subsequently, a retraction step is applied to project the new point back onto the manifold $\mathcal{S}$:
$\footnotesize\theta_{i}^{(t+1)}=\frac{\theta_{i}^{(t)}+\alpha u_{i}^{(t)}}{|\theta_{i}^{(t)}+\alpha u_{i}^{(t)}|},\quad\forall i\in\{1,\dots,LM\}.$ (5)
where $\alpha$ is the step size determined via Armijo backtracking line search [18]. The RCG algorithm iteratively refines the phase shift vector $\boldsymbol{\theta}$ until convergence, ensuring that the unit-modulus constraint (5b) is satisfied at every step.
III-C Computational Complexity Analysis
The proposed algorithm alternates between SCA-based power allocation and RCG-based SIM beamforming. The power allocation subproblem involves solving a convex program with $\mathcal{O}(K^{3})$ complexity per SCA iteration, converging in $t_{1}$ steps [19]. The SIM beamforming subproblem optimizes a length-$LM$ phase vector shift using RCG, with complexity $\mathcal{O}((KLM)^{1.5})$, converging in $t_{2}$ iterations [18]. Thus, the total complexity over $t$ AO iterations is $\mathcal{O}(t(t_{1}K^{3}+t_{2}(KLM)^{1.5}))$.
@@ IV Simulation Results
In this section, numerical simulations are provided to evaluate the performance of the proposed SIM-RSMA system. The BS, positioned at a height of $d_{BS}\!=\!10$ meters, is equipped with N antennas arranged along the x-axis. It utilizes a SIM to perform wave-domain beamforming. The SIM’s layered structure lies parallel to the x-y plane, with all antennas or meta-atoms aligned along the z-axis, ensuring coordinated beam control across the aperture. Moreover, $K$ users are evenly spaced on the y-axis by $d_{u}$ of 10 m as illustrated in Fig. 2. The system operates at a carrier frequency of 28 GHz. The SIM consists of $L$ metasurface layers, each with $M$ meta-atoms arranged in a square grid of size $\sqrt{M}\times\sqrt{M}$. The spacing between adjacent meta-atoms within each layer is set to $\lambda/2$, and the total SIM thickness is $5\lambda$, with an inter-layer spacing of $5\lambda/L$. The noise power at each user is $\sigma_{k}^{2}\!=\!-100\,\text{dBm}$, and the distance-dependent path loss is modeled as $\zeta_{k}\!=\!\beta_{0}(d_{k}/d_{0})^{-\alpha}$, where $\beta_{0}\!=\!-60\,\text{dB}$, $d_{0}\!=\!1\,\text{m}$, and $\alpha\!=\!3.5$ [7]. Unless otherwise mentioned, the parameters used in the simulation are as follows: the number of SIM elements per layer is $M\!=\!64$, the number of users $K\!=\!3$, and the power budget $P_{\max}=30$dBm.
@@
We evaluate the proposed SIM-RSMA scheme against five benchmark schemes: (1) SIM-SDMA that applies private message transmission only, with optimized power allocation and SIM beamforming, (2) SIM-NOMA that uses power-domain multiplexing with SIC decoding [11], assisted by SIM-based analog beamforming, (3) RSMA-UPA that applies RSMA with uniform power allocation and optimized SIM phase shifts, (4) SDMA-UPA that applies SDMA with uniform power allocation and optimized SIM beamforming, and (5) RSMA-Random that evaluates RSMA with optimized power allocation and randomly generated SIM phase shifts.
Fig. 5 presents the average minimum rate achieved as a function of transmit power. At $P_{\max}=40\,\mathrm{dBm}$, SIM-RSMA achieves approximately 22% and 105% higher minimum rate than SIM-SDMA and SIM-NOMA, respectively. These gains are attributed to RSMA’s capability to partially decode interference via the common stream, offering more flexible interference management than SDMA, which treats interference as noise, and NOMA, which relies on strict decoding order and user channel disparity. Uniform power allocation in RSMA and SDMA leads to noticeable performance degradation, while RSMA with random SIM beamforming performs worst, highlighting the need for joint power and phase shift optimization.
Fig. 5 shows the average minimum rate versus the number of meta-atoms per SIM layer $M$. At $M\!=\!100$, the proposed SIM-RSMA scheme with $L\!=\!3$ achieves approximately 40% and 68% higher minimum rate than SIM-SDMA and SIM-NOMA, respectively. Also, the performance consistently improves as the number of SIM layers increases. A significant gain is achieved when $L$ increases from 1 to 2, and although the additional improvement from 2 to 3 layers is relatively smaller, it remains beneficial and contributes to overall system enhancement.. These gains result from the additional degrees of freedom offered by adding more layers to the SIM, which enhance spatial transformations and improve wave-domain beamforming. RSMA benefits more from this capability due to its flexible interference management via common and private stream superposition, unlike SDMA and NOMA, which rely on more rigid spatial or power-domain strategies.
Figure 3: The impact of transmit power ($P_{\max}$) on the minimum rate.
Figure 4: The impact of the number of meta-atoms ($M$) on the minimum rate.
Figure 5: The impact of the number of SIM layers ($L$) on the minimum rate.
Fig. 5 presents the average minimum rate as a function of the number of SIM layers $L$. SIM-RSMA scheme achieves approximately 38% and 64% higher minimum rate than SIM-SDMA and SIM-NOMA, respectively, at $L\!=\!5$ and $M\!=\!81$. Increasing the number of layers improves performance due to enhanced spatial transformation capabilities. However, the gain from increasing $L$ from 1 to 2 is noticeably larger than the gain from increasing $L$ from 4 to 5, indicating diminishing returns as the SIM approaches its wave manipulation capacity. Once sufficient angular resolution and spatial separation are achieved, additional layers yield marginal improvements. Moreover, higher minimum rates are observed with larger numbers of meta-atoms per layer, which indicates the inherent trade-off between increasing the number of layers and expanding the aperture size for effective SIM design.
Fig. 6 shows the impact of increasing the number of users $K$ on the average minimum rate. As expected, the minimum rate decreases as $K$ increases due to higher interference and reduced spatial degrees of freedom. At $K\!=\!6$, the proposed SIM-RSMA scheme with $L\!=\!3$, $M\!=\!256$ achieves 29% and 50% higher minimum rate compared to SIM-SDMA and SIM-NOMA, respectively. Increasing the number of meta-atoms improves performance across all schemes, particularly in dense-user regimes. However, with a small metasurface configuration ($M\!=\!64$), SIM-SDMA and SIM-NOMA converge to similar performance. This is attributed to the increased correlation between the user’s channels and the limited available degrees of freedom or spatial resources. With only 64 elements, the metasurface’s ability to effectively differentiate or allocate resources is limited. As a result, both SIM-SDMA and SIM-NOMA may struggle to fully leverage their respective advantages. Additionally, we observe that distributing $M\!=\!256$ meta-atoms over two layers yields higher minimum rates than placing $M\!=\!529$ meta-atoms in a single-layer configuration. This is because the multi-layer configuration enables more effective spatial separation of the users’ signals through successive phase transformations in the wave domain, even with the same number of total elements.
Figure 6: The impact of the number of users ($K$) on the min rate.

### v2

@@ III Joint Power Allocation and Wave-Based Beamforming: Solution to Problem (P1)
Problem (P1) is a non-convex and challenging optimization problem due to the coupled design of the transmit power variables and the SIM phase shifts in (5a), as well as the non-convex unit-modulus constraints in (5b). Therefore, obtaining the global optimal solution to (P1) is intractable. To address this, we propose an efficient algorithm based on alternating optimization to obtain a near-optimal solution for (P1). In particular, the AO framework alternately optimizes the power allocation and SIM phase shift variables. The corresponding power allocation subproblem and the SIM phase shift subproblem are introduced in subsections III-A and III-B, respectively.
@@ III-A Power Allocation
$(\text{P2}):$ $\max_{\mathbf{p},\mathbf{r}_{c},t}\quad t$ (6a)
$\quad\text{s.t.}\quad r_{c,k}+R_{p,k}\geq t,\quad\forall k\in\mathcal{K},$ (6b)
@@
where $t$ is an auxiliary variable introduced to represent the minimum rate across all users and to enable a smooth reformulation of the original max–min objective. However, problem (P2) remains non-convex due to the non-convexity of constraint (5d) and constraint (6b), as both the private rates $R_{p,k}$ and the common rates $R_{c,k}$ are non-concave functions of the power allocation vector $\mathbf{p}$.
@@
$R_{c,k}=\log_{2}(p_{c}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{c}|^{2}+\sum_{j\in\mathcal{K}}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2})-\eta_{c,k},$ (7)
$R_{p,k}=\log_{2}(\sum_{j\in\mathcal{K}}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2})-\eta_{p,k},$ (8)
@@
$\eta_{\delta,k}=\log_{2}\left(\sum_{j\in\mathcal{K}_{\delta}(k)}p_{j}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}+\sigma_{k}^{2}\right)\!,\quad\delta\in\{\mathrm{c},\mathrm{p}\}.$ (9)
Despite this reformulation, $R_{c,k}$ and $R_{p,k}$ remain non-concave due to the non-concavity of the term $\eta_{\delta,k}$. To address this, we approximate equation (9) using first-order Taylor expansions around the solution obtained at iteration t1t-1 as
$\tilde{\eta}_{\delta,k}^{(t)}\!=\!\eta_{\delta,k}^{(t-1)}+\frac{\sum_{j\in\mathcal{K}_{\delta}(k)}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{j}|^{2}(p_{j}-p_{j}^{(t-1)})}{\ln 2\left(\sum_{i\in\mathcal{K}_{\delta}(k)}p_{i}^{(t-1)}|\mathbf{h}_{k}^{H}\mathbf{F}\mathbf{q}_{i}|^{2}+\sigma_{k}^{2}\right)}.$ (10)
Finally, $R_{c,k}$ and $R_{p,k}$ are now concave functions of the power allocation vector $\mathbf{p}$. Consequently, constraints (5d) and (6b) are convex, and the approximate problem at each iteration is convex and can be efficiently solved using convex optimization solvers such as CVX [cvx]. Hence, we employ the SCA method, which iteratively refines the solution by solving a sequence of convex subproblems. At iteration $t$, SCA approximates the non-concave components $\eta_{c,k}$ and $\eta_{p,k}$ with their first-order convex surrogates $\tilde{\eta}_{c,k}^{(t)}$ and $\tilde{\eta}_{p,k}^{(t)}$, evaluated using the solution from the previous iteration following equation (10).
@@ III-B SIM Beamforming Optimization
Given a fixed power allocation $\mathbf{p}$, the SIM beamforming subproblem aims to maximize the minimum rate of all users by optimizing the phase shifts applied across all metasurface layers. The SIM beamforming subproblem is expressed as
$(\text{P3}):$ $\max_{\mathbf{\Theta}_{l},\mathbf{r}}\quad\min_{k\in\mathcal{K}}\{r_{c,k}+R_{p,k}\}$ (11a)
$\quad\text{s.t.}\quad\text{(5b)},\text{(5d)}.$ (11b)
To facilitate the solution to (P3), we collect the diagonal elements of the phase shift matrices $\{\boldsymbol{\Theta}_{\ell}\}_{\ell=1}^{L}$ into a single stacked vector $\boldsymbol{\theta}\in\mathbb{C}^{LM\times 1}$ defined as $\boldsymbol{\theta}=[\theta_{1,1},\dots,\theta_{1,M},\theta_{2,1},\dots,\theta_{L,M}]^{T}$, where each element $\theta_{\ell,m}$ satisfies the unit-modulus constraint (5b).
The non-convex nature of constraint (5b) makes it particularly difficult to handle (P3) using traditional optimization techniques, necessitating more advanced solution methods. To handle the non-convexity of constraint (5b), we reinterpret the feasible set defined by (5b) as a Riemannian manifold[book] $\mathcal{S}$ as follows
$\mathcal{S}=\{\boldsymbol{\theta}\in\mathbb{C}^{LM}:|\theta_{i}|=1,\quad\forall i\in\{1,\dots,LM\},$ (12)
@@
Moreover, to address the non-smoothness of the objective function in (11a), we adopt a log-sum-exp surrogate to obtain a differentiable approximation as follows
$\mathcal{F}(\boldsymbol{\theta})=-\frac{1}{\beta}\log\left(\sum_{k\in\mathcal{K}}e^{-\beta(r_{c,k}+R_{p,k})}\right)\!,$ (13)
where $\beta>0$ is a parameter that controls the approximation tightness and $\{r_{c,k}\}$ are fixed from the BS step.
Next, we account for the decodability constraint (5d). To do so, we introduce the violation measure
$\Gamma(\boldsymbol{\theta})=\sum_{k=1}^{K}r_{c,k}+\frac{1}{\beta}\log\!\left(\sum_{j=1}^{K}e^{-\beta R_{c,j}(\boldsymbol{\theta})}\right),$ (14)
which becomes positive whenever the constraint is not satisfied. This violation is then penalized through the differentiable function
$\mathcal{F}_{\mathrm{pen}}(\boldsymbol{\theta})=-\mu\cdot\frac{1}{\beta}\ln\!\big(1+e^{\beta\Gamma(\boldsymbol{\theta})}\big),$ (15)
where $\mu>0$ determines the strength of enforcement. The penalty term drives the optimization away from infeasible points. While the penalty ensures feasibility, it may cause vanishing gradients once the constraint is satisfied. To maintain useful search directions in the strictly feasible region, we also introduce the common-stream margin
$\Delta(\boldsymbol{\theta})=-\frac{1}{\beta}\log\!\left(\sum_{j=1}^{K}e^{-\beta R_{c,j}(\boldsymbol{\theta})}\right)-\sum_{k=1}^{K}r_{c,k},$ (16)
which quantifies the slack in the decodability constraint. Based on this margin, we define the reward term
$\mathcal{F}_{\mathrm{rew}}(\boldsymbol{\theta})=\eta\cdot\frac{1}{\beta}\ln\!\left(1+e^{\beta\Delta(\boldsymbol{\theta})}\right),$ (17)
with $\eta\ll\mu$ ensuring that feasibility remains the dominant objective. The reward encourages further improvement of the common-stream capacity once feasibility has been achieved.
By combining the fairness surrogate, penalty, and reward, the smooth SIM objective is given as
$J(\boldsymbol{\theta})=\mathcal{F}_{\mathrm{fair}}(\boldsymbol{\theta})+\mathcal{F}_{\mathrm{pen}}(\boldsymbol{\theta})+\mathcal{F}_{\mathrm{rew}}(\boldsymbol{\theta}).$ (18)
In this formulation, the penalty dominates whenever the decodability constraint is violated, while the reward contributes primarily in the feasible region. These two terms therefore act in a complementary fashion, ensuring both strict constraint satisfaction and continual improvement of the minimum achievable user rate.
This smooth reformulation facilitates the development of the RCG algorithm over the manifold $\mathcal{S}$ to optimize the SIM phase shifts. In each iteration of the RCG algorithm, we compute the Euclidean gradient, $\nabla_{\boldsymbol{\theta}}\mathcal{J}$. Because updates based on the Euclidean gradient may not respect the manifold constraint $\mathcal{S}$, we project it onto the tangent space $T_{\boldsymbol{\theta}}\mathcal{S}$ to obtain the Riemannian gradient, expressed as follows
$\mathbf{g}_{\boldsymbol{\theta}}=\nabla_{\boldsymbol{\theta}}\mathcal{J}-\Re\left(\nabla_{\boldsymbol{\theta}}\mathcal{J}\odot\boldsymbol{\theta}^{*}\right)\odot\boldsymbol{\theta},$ (19)
@@
$\mathbf{u}^{(t+1)}=\mathbf{g}^{(t+1)}+\gamma^{(t+1)}\mathbf{u}^{(t)},$ (20)
where $\gamma^{(t+1)}$ is computed using the Polak–Ribière formula [book]. Subsequently, a retraction step is applied to project the new point back onto the manifold $\mathcal{S}$:
$\theta_{i}^{(t+1)}=\frac{\theta_{i}^{(t)}+\alpha u_{i}^{(t)}}{|\theta_{i}^{(t)}+\alpha u_{i}^{(t)}|},\quad\forall i\in\{1,\dots,LM\}.$ (21)
where $\alpha$ is the step size determined via Armijo backtracking line search [book]. The RCG algorithm iteratively refines the phase shift vector $\boldsymbol{\theta}$ until convergence, ensuring that the unit-modulus constraint (5b) is satisfied at every step.
The proposed algorithm alternates between SCA-based power allocation and RCG-based SIM beamforming. The power allocation subproblem involves solving a convex program with $\mathcal{O}(K^{3})$ complexity per SCA iteration, converging in $t_{1}$ steps. The SIM beamforming subproblem optimizes a length-$LM$ phase vector shift using RCG, with complexity $\mathcal{O}((LM)^{2})$, converging in $t_{2}$ iterations [book]. Thus, the total complexity over $t$ AO iterations is $\mathcal{O}(t(t_{1}K^{3}+t_{2}(LM)^{2}))$.
@@ IV Simulation Results
In this section, numerical simulations are provided to evaluate the performance of the proposed SIM-RSMA system. The BS, positioned at a height of $d_{BS}\!=\!5$ m, is equipped with N antennas arranged along the x-axis. The SIM’s layered structure lies parallel to the x-y plane, with all antennas or meta-atoms aligned along the z-axis, ensuring coordinated beam control across the aperture. Moreover, $K$ users are evenly spaced on the y-axis by $d_{u}$ of 5 m as illustrated in Fig. 2. The system operates at a carrier frequency of 28 GHz. The SIM consists of $L$ metasurface layers, each with $M$ meta-atoms of area $A_{t}=\lambda^{2}/4$ arranged in a square grid of size $\sqrt{M}\times\sqrt{M}$. The spacing between adjacent meta-atoms within each layer is set to $\lambda/2$, and the total SIM thickness is $5\lambda$, with an inter-layer spacing of $5\lambda/L$. The wave propagation distance $d_{\ell}^{m,m^{\prime}}$ between the $m^{\prime}$-th meta-atom on layer $(\ell-1)$ and the $m$-th meta-atom on layer $\ell$ is therefore the Euclidean distance based on their relative spatial coordinates. The noise power at each user is $\sigma_{k}^{2}\!=\!-100\,\text{dBm}$, and the distance-dependent path loss is modeled as $\zeta_{k}\!=\!\beta_{0}(d_{k}/d_{0})^{-\alpha}$, where $\beta_{0}\!=\!-60\,\text{dB}$, $d_{0}\!=\!1\,\text{m}$, and $\alpha\!=\!3.5$ [sim_multi_user]. Unless otherwise mentioned, the parameters used in the simulation are as follows: the number of SIM elements per layer is M=64M\!=\!64, the number of users K=3K\!=\!3, and the power budget $P_{\max}=20$dBm.
@@
We evaluate the proposed SIM-RSMA scheme against four benchmark schemes: (1) SIM-SDMA that applies private message transmission only, with optimized power allocation and SIM beamforming, (2) SIM-NOMA that uses power-domain multiplexing with SIC decoding [rsma0], assisted by SIM-based beamforming, (3) RSMA with fully digital beamforming, and (4) SIM-RSMA-RND that evaluates SIM-RSMA scheme with optimized power allocation and randomly phase shifts
Figure 3: The impact of the transmit power $P_{\max}$ on the minimum rate.
Figure 4: The impact of the number of users $K$ on Jain’s fairness index.
Figure 5: The impact of the number of users $K$ on the minimum rate.
Fig. 5 illustrates the impact of the transmit power $P_{\max}$ on the average minimum rate. The proposed SIM-RSMA scheme consistently outperforms SIM-NOMA across all power levels and achieves performance nearly identical to SIM-SDMA. At $P_{\max}=20\,\mathrm{dBm}$, SIM-RSMA achieves approximately 20% higher minimum rate than SIM-NOMA. The fully digital RSMA scheme provides an upper bound but at the cost of significantly higher hardware complexity and energy consumption. The similarity between SIM-RSMA and SIM-SDMA arises because, in the fully analog configuration, the BS is restricted to power allocation and lacks a digital precoder to shape the common-stream beam, while the SIM is constrained by unit-modulus phase shifts that prevent simultaneous alignment of the common signal toward all users. Since the common rate is limited by the weakest user, increasing it becomes disproportionately costly compared to enhancing private links, causing the optimization to naturally prioritize private streams and making SIM-RSMA effectively behave as a SIM-assisted SDMA configuration.
Figure 5 evaluates the impact of varying the number of users $K$ on Jain’s fairness index for different schemes. As shown, SIM-RSMA, SIM-SDMA, and the digital RSMA schemes maintain near-ideal fairness, with their indices around 0.99 as $K$ increases to 6 users, which demonstrates their robustness in ensuring equitable resource allocation. In contrast, the SIM-RSMA scheme with the sum-rate maximization strategy exhibits significantly poor fairness performance. This is due to the inherent conflict between maximizing aggregate throughput and ensuring equitable user rates. On the other hand, SIM-NOMA exhibits a decline in fairness as $K$ grows, which can be attributed to the constraints of power-domain multiplexing in maintaining multi-user fairness. Additionally, the performance of SIM-RSMA-UPA further emphasizes the necessity of optimized power allocation, as uniform power distribution results in significant fairness deterioration. These findings demonstrate that the SIM-RSMA scheme can achieve strong fairness performance without sacrificing scalability, while also providing additional rate improvements over SDMA and NOMA schemes.
Fig. 5 evaluates the impact of the number of users $K$ on the average minimum rate. As $K$ increases, all schemes experience a gradual decline due to stronger multiuser interference and reduced spatial degrees of freedom. SIM-RSMA and SIM-SDMA achieve similar performance for all $K$, while both outperform SIM-NOMA by approximately 20% on average. With a small aperture (M=64M\!=\!64), SIM-SDMA and SIM-NOMA exhibit nearly identical performance, reflecting limited spatial resolution and high channel correlation. Increasing the aperture to M=256M\!=\!256 widens the performance gap, highlighting the importance of spatial resources. Furthermore, distributing M=256M\!=\!256 meta-atoms over multiple layers yields around 40% higher minimum rates than using a single-layer structure, as the multilayer configuration enables more effective spatial separation through successive wave transformations.
Fig. 8 illustrates the convergence behavior of different schemes. All schemes converge rapidly within approximately 4–9 AO iterations, confirming the efficiency of the proposed optimization framework. For the same metasurface configuration, RSMA and SDMA achieve nearly identical convergence and minimum rates, while both outperform NOMA by about 15% on average. Increasing M=49M\!=\!49 to M=81M\!=\!81 improves the minimum rate by 30%, whereas increasing the L=2L\!=\!2 to L=5L\!=\!5 yields an additional 10% gain.
Figs. 8 and 8 examine the impact of the SIM’s key architectural parameters on the average minimum rate: the number of meta-atoms per layer $M$, which defines the aperture size and spatial resolution, and the number of layers $L$, which determines the depth of wave transformations. RSMA and SDMA achieve nearly identical performance across all configurations, both consistently outperforming NOMA. Increasing $M$ provides a substantial improvement due to enhanced spatial focusing, whereas increasing $L$ yields smaller gains through deeper wave transformations. Specifically, increasing the aperture from M=25M\!=\!25 to M=100M\!=\!100 results in a 297% improvement in the minimum rate, while increasing the number of layers from L=1L\!=\!1 to L=5L\!=\!5 yields a 53% gain. These results indicate that once sufficient beamforming capability is achieved, expanding the aperture size $M$ is generally more effective than adding further layers $L$ in efficient interference suppression.
Figure 6: Convergence behavior of the max-min AO algorithm in iterations.
Figure 7: The impact of the number of meta-atoms $M$ on the minimum rate.
Figure 8: The impact of the number of SIM layers $L$ on the minimum rate.